I need to set up an equation/inequality.
The greater of two numbers is 6 more than 4 times the smaller. Their sum is 41. Find the numbers.

Answer by Boreal(15235)

x=smaller number
4x+6=larger
Their sum is 41
x+4x+6=41
5x+6=41
5x=35
x=7
4x+6=34
Those are the numbers, 7 and 34.